OpenStudy (anonymous):

Solve the equation. Then check your answer, Step by step please. 3/m-4 + 1/3(m-4) = 6/m

OpenStudy (anonymous):

i will try ok

OpenStudy (anonymous):

ok thnkz:)

OpenStudy (anonymous):

m=8+73−−√,8−73−−√

OpenStudy (anonymous):

firstly you multiply m both sides ok

OpenStudy (anonymous):

Sorry wrong answer

OpenStudy (jadeishere):

3m−4+13(m−4)=6m We need to perform a multiplication. The following rule is applied: ABC=ACB In our example, the factors in the new numerator are: 1, (m−4), Notice that all non-fraction factors are placed in the numerator. The only factor in the new denominator is:

OpenStudy (anonymous):

you will get this m*(3/m-4+1/3(m-4)=6/m*m

OpenStudy (anonymous):

wow thnk u so i can write it straght on my paper like that @Jadeishere

OpenStudy (jadeishere):

no 43, sorry there's supposed to be a / mark

OpenStudy (jadeishere):

so 3/m - 4 + (m-4) / 3 = 6/m

OpenStudy (anonymous):

is equation is this m^2-16m=9

OpenStudy (jadeishere):

Sorry my laptop was being stupid

OpenStudy (anonymous):

then what is ans monica

OpenStudy (jadeishere):

m ∈ {(√73+8),(−√73+8)} it's kind of hard to get an answer out of the form Ax^2 + bx + c

OpenStudy (jadeishere):

or, in this case, ax^2 - bx = c

OpenStudy (jadeishere):

Man it took a while to find those signs in word >.<

OpenStudy (anonymous):

thnk u

OpenStudy (jadeishere):

Yeah I hoped I even helped besides giving the answer. I hate doing that but I never really know how to explain myself

OpenStudy (anonymous):

lol thannks you still showed me how to do it tho

OpenStudy (jadeishere):

Yeah :P